1) (a) The vapour pressure of ethanol (C2H5OH) at 15 °C was found to be 5866 Pa, and at 40 °C it was measured to be 2133...

1)

(a) The vapour pressure of ethanol (C2H5OH) at 15 °C was found to be 5866 Pa, and at 40 °C it was measured to be 21332 Pa. Calculate the enthalpy of vaporisation for ethanol.

(b) Draw the Lewis structures for both an ethanol molecule and a water molecule then label and annotate your drawings to illustrate what intermolecular interaction(s) may exist between the two different molecules.
Answer #1

- solution Ti = 15°C + 273 = 288k, P= 5866 Pa Tz = 40°c +273 = 313kg P2 =21332 pa R = 8.314 J moet sHwap (ethance) = ? - weLewis structure of ethanol (CH₃ CH₂OH) c→ 47 x2 = 8e 0-6e- = Ge- H léx6 = 6e 200- H H I 4-5-6-8-H & I-U- I-U- Hico H I (octet